OpenStudy (anonymous):

if a=52 degrees b=9 and c=14 find a to the nearest tenth

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Hint: use the law of cosines a^2 = b^2 + c^2 - 2bc*cos(A) a^2 = 9^2 + 14^2 - 29*14*cos(52) a = ???

OpenStudy (anonymous):

-20,835?

jimthompson5910 (jim_thompson5910):

hmm not sure how you're getting that

jimthompson5910 (jim_thompson5910):

it might help to replace cos(52) with 0.61566147532566 since cos(52 degrees) = 0.61566147532566

OpenStudy (anonymous):

ok one minute

OpenStudy (anonymous):

coming out with the same thing

jimthompson5910 (jim_thompson5910):

a^2 = 9^2 + 14^2 - 29*14*cos(52) a^2 = 9^2 + 14^2 - 29*14*0.61566147532566 a^2 = 81 + 196 - 29*14*0.61566147532566 ... ... ... a = ???

OpenStudy (anonymous):

27.0414432

jimthompson5910 (jim_thompson5910):

81 + 196 - 29*14*0.61566147532566 is equal to 27.0414432

jimthompson5910 (jim_thompson5910):

so a^2 = 27.0414432 a = ???

OpenStudy (anonymous):

is it the sqrt of 27.0414432?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

what is that?

OpenStudy (anonymous):

5.20013876738

jimthompson5910 (jim_thompson5910):

good, you got it

jimthompson5910 (jim_thompson5910):

don't forget to round

OpenStudy (anonymous):

thank you so much

jimthompson5910 (jim_thompson5910):

sure thing